a circular loop of radius 20cm carries a current of 10A , an electrons passes the plane of the loop with speed velocity 2.0 x 10 ^6 , the direction of motion makes an angle 30 with the axis of the circle and passes through its centre , find the magnitude of the magnetic force on the electron at the instant it crosses the plane


Grade:9

1 Answers

ROSHAN MUJEEB
3 years ago
HereR=20cm=20×10−2mR=20cm=20×10-2m
I=10A,V=2×106m/sI=10A,V=2×106m/s
θ=30∘θ=30∘
SinceF=e(V→+B→)F=e(V→+B→)
=eVBsinΘ=eVBsinΘ
=1.6×10−19×2×106μoi2Rsin30=1.6×10-19×2×106μoi2Rsin30
=1.6×10−19×2×106=1.6×10-19×2×106
×4π×10−7×102×20×10−2×12×4π×10-7×102×20×10-2×12
=16π×10(−19)N